The phase difference between two points separated by 0.8 m in a wave of frequency 120 Hz is o90. Then the velocity of wave will be
 

Grade:12

1 Answers

swapna
3 years ago

 

Δϕ=90=2aa
Δx=0.8m
we know that Δϕ=α2π×Δx
So λ=Δϕ2π×Δx=π/22π×0.8
λ=3.2m
v=120Hz & v=vλ
velocityv=120×3.2=384m/s